OpenStudy (anonymous):

What is the y-intercept of y = -1/3 x + 3

terenzreignz (terenzreignz):

To get the y-intercept, just let x = 0; @x=0, y = -(1/3)(0) + 3 y = 3, so the y-intercept is at the point (0, 3)

OpenStudy (lgbasallote):

another way is the y = mx + b b is the y-intercept in this problem m = -1/3 and b = 3 therefore 3 is the y-intercept

OpenStudy (lgbasallote):

it's just the opposite side of the glass ^_6

OpenStudy (lgbasallote):

^_^*